# LiveScene: Language Embedding Interactive Radiance Fields for Physical Scene Rendering and Control

### v1

@@ 3.2 Multi-scale Interaction Space Factorization
$\mathcal{L}_{\text{repuls}}=\frac{1}{\left\|\mathbf{c}_{i}-\mathbf{c}_{j}\right\|+\delta}\mathbf{ELU}(s-\left\|(\tilde{\mathbf{y}}_{m}^{i}\odot\tilde{\mathbf{y}}_{m}^{j})(\mathcal{F}_{s}^{i}-\mathcal{F}_{s}^{j})\right\|),\end{aligned}$
where $\tilde{\mathbf{y}}_{m}^{i}$ and $\tilde{\mathbf{y}}_{m}^{j}$ are the groundtruth mask of rays, and $\mathcal{F}_{s}^{i}$ and $\mathcal{F}_{s}^{j}$ are the last-layer features of interaction probability decoder in Figure. 2. In training iteration, we randomly select ray pairs $(\mathbf{r}_{i},\mathbf{r}_{j})$ and apply $\mathcal{L}_{\text{repuls}}$ to enforce the separation of interactive probability features across local deformable spaces. Meanwhile, we incorporate spatial information by weighting $\mathcal{L}_{\text{repuls}}$ with $\frac{1}{\left\|\mathbf{c}_{i}-\mathbf{c}_{j}\right\|+\delta}$, thereby encouraging the model to focus on local sampling conflicts. The centroid of ray pairs $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ can be obtained by ${\textstyle\sum_{i=1}^{i=M}}w_{i}\mathbf{x}_{i}$, where $w_{i}$ and $\mathbf{x}_{i}$ are the weight and 3D position of each sample, and $M$ is the sample number. Additionally, a probability rejection operation is proposed to truncate the low-probability samples if the deformable probability at $\mathbf{p}$ is smaller than threshold $s$: $d=\arg\max_{i}\Theta(\boldsymbol{\kappa},\boldsymbol{\theta}_{s}(\mathbf{x}))_{i}~{}\text{if}~{}\Theta(\boldsymbol{\kappa},\boldsymbol{\theta}_{s}(\mathbf{x}))_{i}\geq s~{}\text{else}-1.$ The rejected sample is considered to be lying beyond the hyperplane and does not belong to the deformable field. As shown in Figure. 4(b), the proposed operations help the model achieve higher rendering quality, demonstrating their effectiveness in alleviating boundary sampling conflicts.
@@
Figure 4: Illustration of a) boundary sampling conflicts, b) rendering quality comparison, and c) interaction-aware language embedding.
@@ 3.3 Interaction-Aware Language Embedding
Language embedding in interactive scenes is complex and storage-intensive, as 3D distillation faces the dual challenges of high-dimensional optimization and interaction scene variation inconsistency, such as the distinct topological structures of a transformer before and after morph in Figure. 4(c), leading to the failure of SAM [30] segmentation or LERF [28] grounding. Leveraging the proposed multi-scale interaction space factorization in 3.2, we efficiently store language features in lightweight planes by indexing them according to maximum probability sampling instead of 3D fields in LERF. For any sampling point $\mathbf{p}$, we project it onto $\mathbf{p_{d}}=\left[\mathbf{x}|\boldsymbol{\kappa}_{\text{d}}\right]$, retrieving a local language feature group by $d$, and perform bilinear interpolation using $\boldsymbol{\kappa}_{\text{d}}$ to obtain a language embedding that adapts to interactive variable changes from surrounding clip features. Our method has a storage complexity of $\mathbf{O}(C\times\alpha\times\mathbf{dim})$, much smaller than language distillation methods like LERF that operate in 3D scenes, where $\mathbf{dim}$ is the dimension of CLIP feature.
@@ 5 Experiment
In this section, we evaluate LiveScene with existing SOTA methods on CoNeRF Synthetic and Controllable [25] datasets, and the proposed OmniSim and InterReal datasets, demonstrating our superiority in scene-level interactive scenarios reconstruction and language grounding. Detailed implementation can be found in the supplemental materials.
We compare LiveScene with the existing 3D static rendering methods [45, 47, 27, 26], 4D deformable methods [54, 50, 49], and controllable scene reconstruction methods [25, 73]. Note that we reimplemented CoGS [73] based on Deformable Gaussian [72] since the official code is unavailable. Additionally, we extended K-Planes [54] from $\mathbf{C}_{4}^{2}$ planes to $\mathbf{C}_{3+\alpha}^{2}$ planes, denoted as MK-Planes, where $\alpha$ represents the number of interactable objects in dynamic scenes. By leveraging the fact that each instance occupies a distinct region, we further compressed the model, denoted as MK-Planes, requiring only $3+3\alpha$ planes. Detailed implementations can be found in the supplemental materials.
@@ 5.1 View Synthesis Quality Comparison
Evaluation on CoNeRF Synthetic and Controllable Datasets. We report the quantitative results on CoNeRF Synthetic and Controllable scenes in Table. 1. LiveScene outperforms all the existing methods in PSNR, SSIM, and LPIPS metrics on CoNeRF Synthetic scenes with a large margin. In particular, LiveScene achieves 43.349, 0.986, and 0.011 in PSNR, SSIM, and LPIPS, respectively, outperforming the second-best method by 9.894, 0.009, and 0.053. On CoNeRF Controllable, LiveScene achieves the best PSNR of 32.782 and comparable SSIM and LPIPS to the SOTA methods. The results demonstrate the effectiveness of LiveScene in modeling object-level interactive scenarios.
@@
Evaluation on OmniSim and InterReal Datasets. OmniSim dataset is categorized into 3 interaction level subsets: #easy, #medium, and #challenging, based on the number of interactive objects in each scene. As shown in Table. 2, LiveScene achieves the best PSNR, SSIM, and LPIPS on all interaction level subsets of OmniSim, with average PSNR, SSIM, and LPIPS of 33.158, 0.962, and 0.074, respectively. Notably, substantial performance degradation is observed across all methods as the quantity and complexity of interactive objects increase, e.g., CoGS [73] experiences a 3.641 dB PSNR drop from #easy to #challenging. While LiveScene maintains a relatively stable high performance across all subsets, demonstrating its robustness in modeling complex interactive scenarios.
@@
We also divided InterReal dataset into 2 subsets: #medium and #challenging. In  Table. 3, CoGS [73] underperforms compared to LiveScene on the #medium subset and fails to converge when faced with long camera trajectories and a large number of interactive objects in the scene (#challenging), highlighting the limitation of existing methods in modeling real-world interactive scenarios. In contrast, LiveScene achieves the highest PSNR of 28.436 and the lowest LPIPS of 0.185 on the #challenging subset, indicating its superiority in modeling real-world large-scale interactive scenarios.
Figure. 6 presents the novel view synthesis results of LiveScene and the SOTA methods on OmniSim dataset. The results reveal that LiveScene generates more detailed results than SOTA methods, particularly in complex interactive scenarios. For instance, on the #pomaria scene featuring an openable dishwasher, CoNeRF [25] fails to capture details, while CoGS [73] and MK-Planes exhibit residual artifacts. In contrast, our method accurately reconstructs the internal details. Another challenge arises in the #rs scene, where other methods struggle to reconstruct distant and static objects. In comparison, our method not only overcomes the challenging problem of dramatic topology changes in interactive scenes but also maintains the ability of high-quality static scene reconstruction.
@@
Depth L1 Error $\downarrow$
scene CoNeRF MK-Planes CoGS LiveScene
#seq04 0.029 0.037 0.381 0.018
#seq08 0.018 0.301 0.655 0.025
#seq14 0.042 0.103 0.810 0.039
#seq15 0.019 0.568 0.690 0.021
#seq17 0.282 0.706 0.019
#seq19 0.136 0.689 0.034
#avg 0.027 0.238 0.655 0.026
Figure 7: Structure Reconstruction Performance on OmniSim Dataset. left): Our method surpasses most previous works on chosen subsets. right): LiveScene excels in both reconstructing intricate details of interactive objects and preserving the geometric integrity of static background.
5.2 Interactive Scenes Geometry Comparison
To evaluate the completeness of the topological structure of interactive objects, we employ the depth L1 error metric. As shown in Figure. 7, our method outperforms SOTA methods on scenes from OmniSim. While existing methods excel in RGB image rendering, they struggle with depth structure representation. Specifically, CoNeRF [25] performs relatively well in #seq08 and #seq15 but fails in large-scale scenes (#seq17 and #seq19). CoGS [73] exhibit notable artifacts in the depth map. Moreover, MK-Planes also fails to recover depth around interactive objects. In contrast, our method achieves the lowest depth error and renders satisfying depth maps, demonstrating its accurate interactive scene modeling capabilities.
@@ 5.3 Language Grounding Comparison
We assess the language grounding performance on OmniSim dataset using mIOU metric. Figure. 8 suggests that our method obtains the highest mIOU score, with an average of 86.86. In contrast, traditional methods like LERF [28] encounter difficulties in locating objects precisely, with an average mIOU of 21.74. Meanwhile, 2D methods like SAM [30] fail to accurately segment the whole target under specific viewing angles, as objects appear discontinuous in the image. Conversely, our method perceives the completeness of the object and has clear knowledge of its boundaries, demonstrating its advantage in language grounding tasks.
@@ 5.4 Ablation Study
In this section, we present ablative studies to investigate the effectiveness of each component in LiveScene. We selected 1 scene from the #medium subset and 2 scenes from the #easy subset for OmniSim dataset, and 1 scene each from the #medium and #challenging settings for InterReal. Notably, ground-truth state quantities are only available in OmniSim, but not in InterReal. Therefore, we use GT quantities on OmniSim and introduce a learnable variable on InterReal to infer state changes. Figure. 9 reports the rendering quality and grounding performance for #1 and #6.
As illustrated in Table. 4, the multi-scale factorization significantly improves the rendering performance on both datasets, with PSNR on OmniSim increasing from 31.74 to 35.094, shown in #1. Introducing learnable variables for each frame (#2) yields corresponding improvements on InterReal dataset since this latent code can perceive the change in object states. The feature repulsion loss and probability rejection (#3 and #4) together make rendering quality better in InterReal as well as in OmniSim dataset. As for grounding, #5 shows that rendering embeddings along a ray [31, 79] struggles to locate objects precisely. Ensuring view consistency further boosts grounding performance on OmniSim, as demonstrated in #6.

### v2

@@ 3.2 Multi-scale Interaction Space Factorization
$\mathcal{L}_{\text{repuls}}=\mathbf{ELU}(K-\left\|(\mathbf{M}_{i}\odot\mathbf{M}_{j})(\mathcal{F}_{i}-\mathcal{F}_{j})\right\|),\end{aligned}$
where $K$ is a constant hyperparameter, $\mathbf{M}_{i}$ and $\mathbf{M}_{j}$ are the ground truth mask of rays. $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ are the last-layer features of interaction probability decoder in Figure. 2. During training, we randomly select ray pairs and apply $\mathcal{L}_{\text{repuls}}$ to enforce the separation of interactive probability features across local deformable spaces. The initiative is inspired by [24], which demonstrated the efficacy of repulsive forces in disambiguating 3D segmentation results.
Additionally, the probability rejection is proposed to truncate the low-probability samples if the maximum deformable probability $\mathbf{P}$ at sample $\mathbf{p}$ is smaller than threshold $s$, and selects the background feature directly. The operation is defined as:
$u=\left\{\begin{array}[]{ll}\operatorname*{arg\,max}_{i}\{\mathbf{P}_{i}\},&\text{if}\quad\mathbf{P}_{i}\geq s\\ -1,&\text{ otherwise }\end{array}\right..$ (5)
As shown in Figure. 4(b), the proposed operations help the model achieve higher rendering quality, demonstrating their effectiveness in alleviating boundary sampling conflicts.
@@
Figure 4: Illustration of a) boundary sampling conflicts, b) rendering quality comparison.
@@ 3.3 Interaction-Aware Language Embedding
Language embedding in interactive scenes is complex and storage-intensive, as 3D distillation faces the dual challenges of high-dimensional optimization and interaction scene variation inconsistency, such as the distinct topological structures of a transformer toy before and after transformation, leading to the failure of SAM [25] segmentation or LERF [23] grounding. As shown in Figure. 2, leveraging the proposed multi-scale interaction space factorization of 3.2, we efficiently store language features in lightweight planes by indexing them according to maximum probability sampling instead of 3D fields in LERF. For any sampling point $\mathbf{p}$, we project it onto $\mathbf{p}_{u}=\left[\mathbf{x}|\boldsymbol{\kappa}(u)\right]$, retrieving a local language feature group by index $d$, and perform bilinear interpolation using $\boldsymbol{\kappa}_{u}$ to obtain a language embedding that adapts to interactive variable changes from surrounding clip features. By interpolating language embeddings, our method not only perceives topological structure changes but also achieves a storage complexity of $\mathbf{O}(C\times\alpha\times\mathbf{dim})$, much smaller than language distillation methods like LERF that operate in 3D scenes, where $\mathbf{dim}$ is the dimension of CLIP feature.
@@ 5 Experiment
Baselines. We compare LiveScene with the existing 3D static rendering methods [38, 39, 22, 21], 4D deformable methods [10, 42, 41], and controllable scene reconstruction methods [20, 63]. Note that we reimplemented CoGS [63] based on Deformable Gaussian [62] since the official code is unavailable. Additionally, we extended K-Planes [10] from $\mathbf{C}_{4}^{2}$ planes to $\mathbf{C}_{3+\alpha}^{2}$ planes, denoted as MK-Planes, where $\alpha$ represents the number of interactable objects in dynamic scenes. By leveraging the fact that each instance occupies a distinct region, we further compressed the model, denoted as MK-Planes, requiring only $3+3\alpha$ planes.
Implementation. LiveScene is implemented in Nerfstudio [51] from scratch. We represent the field as a multi-scale feature plane with resolutions of $512\times 256\times 128$, and feature dimension of 32. The proposal network adopts a coarse-to-fine sampling process, where each sampling step concatenates the position feature and the state quantity as the query for the 4D deformation probability field, which is a 1-layer MLP with 64 neurons and ReLU activation. We use the Adam optimizer with initial learning rates of 0.01 and a cosine decay scheduler with 512 warmup steps for all networks. The model is trained with an NVIDIA A100 GPU for 80k steps on the OmniSim dataset and 100k steps on the InterReal dataset, using a batch size of 4096 rays with 64 samples. More implementation can be found in the supplemental materials.
@@ 5.1 View Synthesis Quality Comparison
Evaluation on CoNeRF Synthetic and Controllable Datasets. We report the quantitative results on CoNeRF Synthetic and Controllable scenes in Table. 1. LiveScene outperforms all the existing PSNR, SSIM, and LPIPS metrics methods on CoNeRF Synthetic scenes with a large margin. In particular, LiveScene achieves 43.349, 0.986, and 0.011 in PSNR, SSIM, and LPIPS, respectively, outperforming the second-best method by 9.894, 0.009, and 0.053. On CoNeRF Controllable, LiveScene achieves the best PSNR of 32.782 and comparable SSIM and LPIPS to the SOTA methods. According to the novel view synthesis results in Figure. 6, LiveScene achieves more detailed and higher rendering quality, demonstrating the effectiveness of LiveScene in modeling object-level interactive scenarios.
@@
Evaluation on OmniSim Datasets. OmniSim dataset is categorized into 3 interaction level subsets: #easy, #medium, and #challenging, based on the number of interactive objects in each scene. As shown in Table. 2, LiveScene achieves the best PSNR, SSIM, and LPIPS on all interaction level subsets of OmniSim, with average PSNR, SSIM, and LPIPS of 33.158, 0.962, and 0.074, respectively. Notably, substantial performance degradation is observed across all methods as the quantity and complexity of interactive objects increase, e.g., CoGS [63] experiences a 3.641 dB PSNR drop from #easy to #challenging. While LiveScene maintains a relatively stable high performance across all subsets, demonstrating its robustness in modeling complex interactive scenarios.
Evaluation on InterReal Datasets. We divided InterReal dataset into #medium and #challenging subsets. In  Table. 3, CoGS [63] underperforms compared to LiveScene on the #medium subset and fails to converge when faced with long camera trajectories and a large number of interactive objects in the scene (#challenging), highlighting the limitation of existing methods in modeling real-world interactive scenarios. In contrast, LiveScene achieves the highest PSNR of 28.436 and the lowest LPIPS of 0.185 on the #challenging subset, indicating its superiority in modeling real-world large-scale interactive scenarios.
Figure 6: View Synthesis Visualization on CoNeRF Controllable Dataset. The proposed method achieves higher-quality rendering results compared with the existing methods.
@@
View Synthesis Visulization. Figure. 7 presents the novel view synthesis results of LiveScene and the SOTA methods on OmniSim dataset. The results reveal that LiveScene generates more detailed results than SOTA methods, particularly in complex interactive scenarios. For instance, on the #pomaria scene featuring an openable dishwasher, CoNeRF [20] fails to capture details, while CoGS [63] and MK-Planes exhibit residual artifacts. In contrast, our method accurately reconstructs the internal details. Another challenge arises in the #rs scene, where other methods struggle to reconstruct distant and static objects. In comparison, our method not only overcomes the challenging problem of dramatic topology changes in interactive scenes but also maintains the ability to reconstruct high-quality static scenes.
@@ 5.2 Language Grounding Comparison
We assess the language grounding performance on the OmniSim dataset using mIOU metric. Figure. 8 suggests that our method obtains the highest mIOU score, with an average of 86.86. In contrast, traditional methods like LERF [23] encounter difficulties in locating objects precisely, with an average mIOU of 21.74. Meanwhile, 2D methods like SAM [25] fail to accurately segment the whole target under specific viewing angles, as objects appear discontinuous in the image. Conversely, our method perceives the completeness of the object and has clear knowledge of its boundaries, demonstrating its advantage in language grounding tasks.
@@
Figure 10: Learning process of the probability fields from 0 to 1000 training steps. The model progressively converges to the vicinity of the interactive objects, establishing interactive regions.
@@ 5.3 Ablation Study
In this section, we present ablative studies to investigate the effectiveness of each component in LiveScene. We selected 1 scene from the #medium subset, 2 scenes from the #easy subset of OmniSim dataset, and 1 scene each from the #medium and #challenging settings for InterReal. Notably, ground-truth state quantities are only available in OmniSim, not in InterReal. Therefore, we use GT quantities on OmniSim and introduce a learnable variable on InterReal to infer state changes. Figure. 9 reports the rendering quality and grounding performance for #1 and #6.
Effectiveness of components. As illustrated in Table. 4, the multi-scale factorization significantly improves the rendering performance on both datasets, with PSNR on OmniSim increasing from 31.74 to 35.094, shown in #1. Introducing learnable variables for each frame (#2) yields corresponding improvements on InterReal dataset since this latent code can perceive the change in object states. The feature repulsion loss and probability rejection (#3 and #4) together make rendering quality better in InterReal as well as in OmniSim dataset. As for grounding, #5 shows that rendering embeddings along a ray [26, 69] struggles to locate objects precisely. Ensuring view consistency further boosts grounding performance on OmniSim, as demonstrated in #6.
Probability field training. We provide additional experiments in Figure. 10 of the disjoint regions to illustrate the learning process of the probability field from 0 to 1000 training steps. The results demonstrate a clear trend that, as training advances, the proposed method can progressively converge to the vicinity of the interactive objects, thereby establishing interactive regions. With the establishment of the probability field, the model can focus on different interactive objects and guide the sampling process by maximizing probability, thereby achieving disentanglement of interactive scenes.